1

Base Year Fruit Apples Bananas 7,000 bunches Oranges Quantity Price $2 per bag $3 per bunch $5 per bag 4,000 bags 7,000 bags Current Year Fruit Quantity Apples Bananas 15.000 bunches Oranges Price $3 per bag $2 per bunch $7 per bag 5,000 bags 28,000 bags Consider an economy that produces only three types of fruit: apples bananas, and oranges. In the base year (a few years ago), the production and price data are listed in the tables to the right. In the base year, nominal GDP was $ In the current year, nominal GDP is $ (Round both answers to the nearest whole number.) The percentage increase in nominal GDP since the base year is (Enter your response as a percentage rounded to one decimal place.) In the base year, real GDP was S In the current year, real GDP is (Round both answers to the nearest whole number.) The percentage increase in real GDP since the base year is %. Enter your response as a percentage rounded to one decimal place.) The GDP deflator for the base year is The GDP deflator for the current year is (Round both answers to one decimal place.) The inflation rate (which equals the percentage increase in the GDP deflator) since the base year is 1 %. (Enter your response as a percentage rounded to one decimal place.)
Answer #1

Nominal GDP = Price in the given year * Quantity in the given year

Therefore in the base year ,

nominal GDP =$2* 4000 + $3*7000 + $5*7000

= $64000

Similarly in the current year ,

nominal GDP = $3* 5000 + $2*15000 + $7*28000

= $241000

Percentage increase in nominal GDP =

{(nominal GDP in current year - nominal GDP in base year )/ nominal GDP in base year}

Therefore, percentage increase since the base year

= { ($241000 - $64000)/$64000} *100 = 276 %

Real GDP = Price in the base year * Quantity in the given year

The real and nominal GDP would be equal to each other in the base year. However, for the current year they would be different (It is evident from the given equation) .

real GDP for current year =$2* 5000 + $3*15000 + $5*28000 = $195000

Percentage increase in real GDP = { (real GDP in current year - real GDP in base year)/ real GDP in base year} *100

= {($195000 - $64000) $64000} *100 = 204%

GDP deflator = (Nominal GDP / Real GDP)*100

For base year GDP deflator = ($64000/$64000)*100 = 100

Similarly, for current year GDP deflator

= ($241000/$195000) *100 = 123.58

Inflation = percentage increase in GDP deflator

= {( 123.58 - 100) /100} *100 = 23 percentage